Spin Cycle Architecture uses three activity pools to apply overhead to its projects. Each activity has a cost driver used to allocate the overhead costs to the projects. The activities and related overhead costs are as follows: initial concept formation $41,000; design $334,000; and construction oversight $116,550. The cost drivers and estimated use are as follows.

Activities

Cost Drivers

Estimated Use of Cost
Drivers per Activity

Initial concept formation Number of project changes 25
Design Square feet 167,000
Construction oversight Number of months 105


(a) Compute the predetermined overhead rate for each activity.

Overhead Rate

Initial concept formation $ per project change
Design $ per square foot
Construction oversight $ Per month


(b) Classify each of these activities as unit-level, batch-level, product-level, or facility-level.

Activities

Level

Initial concept formation

Product-levelUnit-levelFacility-levelBatch-level

Design

Batch-levelProduct-levelFacility-levelUnit-level

Construction oversight

Batch-levelFacility-levelUnit-levelProduct-level
Answer #1

Solution-

a)

Overhead Rate
Initial concept formation $1640 Per project change
Design $2 per square foot
Construction oversight $1110 per month

Working- Predetermined overhead rate = Estimated overhead / Estimated allocation base

*Initial concept formation = ($41000/25) = $1640

* Design = ($334000 / 167000) = $2

* Construction oversight = ($116550/105) = $1110

b)

Activities Level
Initial concept formation Product level
Design Unit level
Construction oversight Batch level